OpenStudy (anonymous):

A loan of R90 000, 00 is taken out and repaid by equal monthly instalments over a 3 year period. Interest is set at 7, 5% p.a. compounded monthly. Calculate the monthly repayment.

OpenStudy (anonymous):

formula: (Co x i)/ 1-(1+i)^-n 90.000 x 0,075/12 / 1-(1+0,075/12)^-36 = 562,5/ 0,2009 = 2.799 monthly instalment